OpenStudy (anonymous):

Selena buys several shares of stock at $32 per share and sells them at $28.32 per share. What is the percent of decrease in the price of the stock? (Points : 1) 12.0% 11.5% 11.0% 10.5%

OpenStudy (some.random.cool.kid):

$32.00 - $28.32 = $3.68 68/32 = .115 = 11.5%

OpenStudy (anonymous):

32-28.32=3.86 3.86/32.115 .115*100=11.5%

OpenStudy (some.random.cool.kid):

thats how you solve it

OpenStudy (some.random.cool.kid):

lol you wrote the same thing i wrote i guess were in agreement